OpenStudy (sleepyjess):

Find an equation in standard form for the hyperbola with vertices at (0, ±10) and asymptotes at \(y =\pm \dfrac 56 x\)

OpenStudy (sleepyjess):

I think it is \(\dfrac{x^2}{100} - \dfrac{y^2}{144} = 1\)

OpenStudy (sleepyjess):

@ganeshie8

OpenStudy (rational):

give vertices are \(0,~\pm 10\) theat means the hyperbola opens up and down, yes ? (Vertical hyperbola)

OpenStudy (sleepyjess):

Yes

OpenStudy (sleepyjess):

oh, so it would be \(\dfrac{y^2}{100} - \dfrac{x^2}{144} = 1\)

OpenStudy (sleepyjess):

Thanks! :)

OpenStudy (rational):

np:)